OpenStudy (anonymous):

I'll give a medal and become a fan of anyone who can help me solve this: "The Richter scale is used to measure the magnitude of an earthquake. The magnitude, R, is given by the equation R = 0.65log(0.39E) + 1.45, where E is the energy (measured in kilowatt-hours) released by the earthquake. Determine the amount of energy released by an earthquake of magnitude 7.5"

OpenStudy (anonymous):

I already plugged in 7.5, so 7.5 = 0.65log(0.39E) + 1.45

OpenStudy (anonymous):

I just don't know how to solve for E from there.

OpenStudy (anonymous):

Oh, I guess I could move the 1.45. 6.05 = 0.65log(0.39E)

OpenStudy (anonymous):

@phi ? Again I'm sorry to keep tagging you ._. I try to wait until someone else comes to help but usually nobody does.

OpenStudy (phi):

there is one other "obvious" step. how do you get the log() by itself on the right side ?

OpenStudy (anonymous):

hmm.. divide both sides by 0.39E?

OpenStudy (anonymous):

oh wait...

OpenStudy (anonymous):

divide by 0.65?

OpenStudy (anonymous):

I'm not quite sure...

OpenStudy (phi):

the way to remember (or figure it out) is if you have 2x = 5 if you divide 2 by itself you get 1 so divide both sides by 2 x = 5/2 in this case 0.65log(0.39E) = 6.05 and you want to make the "0.65 go away"

OpenStudy (phi):

0.65 divided by itself is 1

OpenStudy (anonymous):

oh okay so then it would end up being log(0.39E) = 6.05/0.65 or log(0.39E) = 9.308 (rounded)

OpenStudy (anonymous):

right?

OpenStudy (phi):

yes, that is good. next, we get rid of the log. make the log the exponent of the base this is hard to say but we do it this way: log(a) (base 10) is "undone" by 10^log(a) which gives a

OpenStudy (phi):

if we start with log(0.39E) = 9.308 make each side the exponent of the base of the log (10 is assumed) 10^(log(0.39E) ) = 10^9.308 that ugly stuff on the left side is 0.39E (we undid the log) 0.39 E = 10^9.308

OpenStudy (phi):

solve for E by dividing by 0.39

OpenStudy (anonymous):

oh okay (: hold on I'll post my answer once I have it

OpenStudy (anonymous):

For my final answer I got E = 5,211,171,822.913

OpenStudy (phi):

I would round that to E = 5.21 * 10^9 kw-hours in scientific notation.